Dr. David Bryan Carpenter Flat 8, Broughton House 28/30 Bullar Road Bitterne Park Southampton, S018 1GS

December 12, 1995

The Visa Service, 2 Northdown Street, Kings Cross, London, N1 9BG

Dear Visa Service,

Further to our previous correspondence I enclose

• My passport.

• Completed application form for US visa.

• 3 original copies of an IAP-66 form.

• Letter from future employer (Syracuse University).

• 1 colour photograph

• Cheque for £24.90 made payable to “The Visa Service”.

• Your receipt for my deposit

Yours sincerely,